OpenStudy (anonymous):

Malin and Shawn begin with the same number, t. Malin subtracts 6 from t and then divides the result by 6. Shawn adds 7 to t and then divides the result by 7. If Malin's final answer is the same as Shawn's final answer, what was the number t?

OpenStudy (anonymous):

Medal for who helps.

OpenStudy (anonymous):

First write it out: Malin \[\frac{ (t-6) }{ 6 }\] and

OpenStudy (anonymous):

Shawn \[\frac{ (t+7) }{ 7 }\]

OpenStudy (anonymous):

Oh I think I can solve it.

OpenStudy (anonymous):

ok go ahead and post your answer when your done and i can double check for you if you like

OpenStudy (anonymous):

t=84? i'm not sure

OpenStudy (anonymous):

if you think its 84 solve it normally.

OpenStudy (anonymous):

oh i's not T_T

OpenStudy (anonymous):

I have to go.

OpenStudy (anonymous):

set it up like a proportion

OpenStudy (anonymous):

it is 84 (t - 6)/6 = (t + 7)/7 (84 - 6)/6 = (84 + 7)/7 78/6 = 91/7 13 = 13 (correct)